# Guided Diffusion Sampling on Function Spaces with Applications to PDEs

### v1

@@ G.6 Additional experiments on Navier-Stokes with boundary conditions
G.7 Design Choices of DiffusionPDE and FunDPS
In order to show the effectiveness of our design choice, we include ablation study on Darcy Flow PDE by adding different components to achieve FunDPS starting from DiffusionPDE. Details are shown in Table 9, and FunDPS’s superiority can be observed upon adding the components.
@@ G.7 Design Choices of DiffusionPDE and FunDPS
Table 9: Ablation study on the design choices of DiffusionPDE and FunDPS, tested on forward and inverse Darcy Flow problems. Functional diffusion noise significantly boosts capability. FunDPS achieves superior performance with fewer steps by incorporating both functional noise and neural operators.
Appendix H Qualitative Comparison
Here we conduct qualitative comparisons between the prediction of FunDPS and DiffusionPDE on Darcy Flow (Figure 9), Poisson (Figure 10), Helmholtz (Figure 11) and Navier-Stokes (Figures 12 and 13) problems.
@@ Appendix H Qualitative Comparison
Appendix I Reproducibility of DiffusionPDE
We replicated DiffusionPDE’s results using their provided code and weights. Due to reproducibility issues, we reran all experiments in communication with DiffusionPDE’s authors and present the comparison in Table 10.
Darcy Flow Poisson Helmholtz Navier-Stokes Navier-Stokes (BCs)
Appendix J Impact Statement

### v2

@@ H.7 Additional experiments on Navier-Stokes with boundary conditions
H.8 Design Choices of DiffusionPDE and FunDPS
In order to show the effectiveness of our design choice, we include an ablation study on Darcy Flow PDE by adding different components to achieve FunDPS starting from DiffusionPDE. Details are shown in Table˜12, and FunDPS’s superiority can be observed upon adding the components.
Table 12: Ablation study on the design choices of DiffusionPDE and FunDPS, tested on forward and inverse Darcy Flow problems. Functional diffusion noise significantly boosts capability. FunDPS achieves superior performance with fewer steps by using functional noise and neural operators.
@@ H.8 Design Choices of DiffusionPDE and FunDPS
Appendix I Reproducibility of DiffusionPDE
We replicated DiffusionPDE’s results using their provided code and weights. Due to reproducibility issues, we reran all experiments in communication with DiffusionPDE’s authors and present the comparison in Table˜13.
Darcy Flow Poisson Helmholtz Navier-Stokes Navier-Stokes (BCs)
Appendix J Impact Statement
Appendix K Qualitative Comparison
Here we conduct qualitative comparisons between the prediction of FunDPS and DiffusionPDE on Darcy Flow (Figure˜8), Poisson (Figure˜9), Helmholtz (Figure˜10) and Navier-Stokes (Figures˜11 and 12) problems.